OpenStudy (anonymous):

y= sinx [pi/6-pi/9]. Is it 1 to 1 y= sinx [pi/6-pi/9]. Is it 1 to 1 @Mathematics

OpenStudy (whynot):

yes

OpenStudy (anonymous):

what does \[[\frac{\pi}{6}-\frac{\pi}{9}]\] mean?